1. Department G had 2,040 units 25% completed at the beginning of the period, 13,200 units...

1.

Department G had 2,040 units 25% completed at the beginning of the period, 13,200 units were completed during the period, 1,700 units were 20% completed at the end of the period, and the following manufacturing costs debited to the departmental work in process account during the period:

Work in process, beginning of period $28,300
Costs added during period:
  Direct materials (12,860 units at $8) 102,880
  Direct labor 88,200
  Factory overhead 29,400

All direct materials are placed in process at the beginning of production and the first-in, first-out method of inventory costing is used. The total cost of 2,040 units of beginning inventory which were completed during the period (round unit cost calculations to four decimal places and round your final answer to the nearest dollar)?

2.

Ramapo Company produces two products, Blinks and Dinks. They are manufactured in two departments, Fabrication and Assembly. Data for the products and departments are listed below.


Product
Number of
Units
Direct Labor Hours
Per Unit
Machine Hours
Per Unit
Blinks 1,109 3 8
Dinks 1,895 5 9

All of the machine hours take place in the Fabrication department, which has an estimated overhead of $117,500. All of the labor hours take place in the Assembly department, which has an estimated total overhead of $85,100.

Ramapo Company uses a single plantwide overhead rate to apply all factory overhead costs based on direct labor hours. The factory overhead allocated per unit of Dinks is

a.$63.32

b.$255.33

c.$79.15

d.$17.92

3.

Reynold's Company has a product with fixed costs of $305,000, a unit selling price of $26, and unit variable costs of $19. The break-even sales (units) if the variable costs are decreased by $3 is

a.43,571 units

b.19,063 units

c.30,500 units

d.16,053 units
Answer #1
1
Direct materials Conversion costs
To complete beginning work in process 0 1530 =2040*(1-25%)
Units started and completed 11160 11160 =13200-2040
Work in process, ending 1700 340 =1700*20%
Equivalent units of production 12860 13030
Direct materials Conversion costs
Costs added during period 102880 117600
Divide by Equivalent units of production 12860 13030
Cost per Equivalent unit 8.0000 9.0253
Direct materials Conversion costs Total
Work in process, beginning 0 0 28300
To cmplete Work in process, beginning 0 13809 13809 =1530*9.0253
Total cost of 2,040 units of beginning inventory 42109
Total cost of 2,040 units of beginning inventory 42109
2
Total overhead 202600 =117500+85100
Divide by Total Direct labor hours 12802 =(1109*3)+(1895*5)
Single plantwide overhead rate 15.83
Factory overhead allocated per unit of Dinks 79.15 =5*15.83
Option C $79.15 is correct
3
Fixed costs 305000
Divide by Unit Contribution margin 10 =26-(19-3)
Break-even sales (units) 30500
Option C 30,500 units is correct